OpenStudy (maddss):

Find the coordinates of the foci of the hyperbola y^2/9-x^2/16=1. A) (4, 0) and (-4, 0) B) (5, 0) and (-5, 0) C) (0, square root 7) and (0, -square root 7) D) (0, 5) and (0, -5)

OpenStudy (anonymous):

-9=-16+c^2 so c^2=7 c=sqrt7 or c=-sqrt7 so answer must be C.

Directrix (directrix):

I did not get C as the correct option. Perhaps I made an error.

OpenStudy (anonymous):

definetely you made an error :)

Directrix (directrix):

a^2 = 9 and b^2 = 16 c^2 - a^2 = b^2 c^2 = 9 + 16 c^2 = 25 c = +5 or -5 Foci at (0, -5) and (0, 5).

OpenStudy (maddss):

So what's the answer?

OpenStudy (goformit100):

C) (0, square root 7) and (0, -square root 7)

OpenStudy (maddss):

thanks :)

Directrix (directrix):

@maddss Take a look at the graph and the statement of the foci given and see what you think. http://www.wolframalpha.com/input/?i=y%5E2%2F9+-+x%5E2%2F16+%3D+1+foci

OpenStudy (anonymous):

@marsss is wrong. 16+9=25, sqrt of 25 is 5, and the graph is up and down, so the foci would be (0,5)(0,-5)